Working together, it takes two different sized hoses 30 minutes to fill a small swimming pool. If it takes 40 minutes for the larger hose to fill the swimming pool by itself, how long will it take the smaller hose to fill the pool on its own?
Answer #1

time taken to fill when working together = 30 min
So, in 1 min, part of pool filled = 1/30

Larger hose fill the tank in 40 min
So,
part of tank filled in 1 min = 1/40

Let time taken by smaller hose to fill tank = x minutes
So,
part of tank filled in 1 min = 1/x

part of tank filled in 1 min if both works together = 1/40 + 1/x

So,
1/40 + 1/x = 1/30
1/x = 1/30 - 1/40
1/x = (4-3)/120
1/x = 1/120
x = 120 minutes

Answer: 120 minutes